OpenStudy (anonymous):

expand (1+1)^n by using binomial theorem and then simplify

OpenStudy (anonymous):

I think I already answered this: \[n!/n! + n!/((n-1)!(1!))+n!/((n-2!)(2!))+...+(n!)/((1!)(n-1)!)+ n!/n!\]

OpenStudy (anonymous):

here \[t _{r}=C _{n}^{r}\] \[s_{r}=\sum_{r=0}^{n}t _{r}\] nC0+nC1+nC2+....nCn=2^n. so simplifying you get 2^n. this is quiet obvious from the above relation that (1+1)^n=2^n